OpenStudy (anonymous):

Precalc help please? Write the equation of the hyperbola whose center is at the origin and has a vertical transverse axis. The equations of the asymptotes are 6x+2y=0 and 6x-2y=0. So, I keep getting (y^2/9)-(x^2/4)=1, but that can't be it because the asymptotes of this are +/-3/2x and they're supposed to be +/-3x?

OpenStudy (anonymous):

@ganeshie8 can you help me with this?

OpenStudy (beccaboo333):

I'm sure he can :3 he's like amazing with math <3

ganeshie8 (ganeshie8):

lol no, you're being kind of me becca :) conics are tricky

ganeshie8 (ganeshie8):

(y^2/9)-x^2=1

ganeshie8 (ganeshie8):

this wud give u the required asymptotes

ganeshie8 (ganeshie8):

may i knw how u worked the equation you're having ? :)

OpenStudy (anonymous):

Oh okay, could you explain how to get that?

ganeshie8 (ganeshie8):

i have just fixed it from the asymptotes, but im sure there must be a standard method

OpenStudy (anonymous):

And as for what I did before, since b^2 is c^2-a^2 I think I just got the value of c messed up.

ganeshie8 (ganeshie8):

you said you're getting 2 extra, so i chopped off 2^2 from the denominator of x^2 so that asymptotes look correct

ganeshie8 (ganeshie8):

do u work it by drawing rectangle ?

OpenStudy (anonymous):

Oh okay, I was basically just guessing and checking before actually :/

OpenStudy (anonymous):

I have to go, but thank you for your help even though I still don't fully understand :P

OpenStudy (beccaboo333):

Sorreh I didn't help. But Ganesh is well known for helping in math :3

ganeshie8 (ganeshie8):

with the given asymptotes we can oly find the relation between "a" and "b" \(\large \frac{a}{b} = \pm 3\)

ganeshie8 (ganeshie8):

to find the exact equation of hyperbola, we will need atleast "one point" on the hyperbola. since we're just given asymptotes and not given any point, there will be infinitely many hyperbolas with these asymptotes

ganeshie8 (ganeshie8):

for ex, below is another hyperbola having the same asymptotes :- \(\large \frac{y^2}{18} - \frac{x^2}{2} = 1\)

ganeshie8 (ganeshie8):

basically, asymptotes are not sufficient to uniquely identify a hyperbola. we need a "point" also

ganeshie8 (ganeshie8):

when u get back, let me knw if smthng doesnt make sense :)

OpenStudy (anonymous):

Yeah, I think I understand :) So the fact that the center is at (0, 0) doesn't help us since that isn't a point on the actual hyperbola itself?